A frictionless piston-cylinder device contains 8 kg of superheated water vapor at 500 kPa and 300oC. Steam is now cooled at constant pressure until 70 percent of it, by mass condenses. Determine the work done and heat transfer during this process.


1
Expert's answer
2021-05-31T06:18:27-0400

The superheated steam at pressure, "P_1=500kPa"

Temperature, "T_1= 300^0 C"

"mass, m=8 kg"

Therefore, The enthalpy at initial state= "H_1=m*h_1=8*(3064.6)=24516.8 kJ"

If it is cooled at the same pressure, steam becomes saturated at temp "T_2=152^0 C".

At "P_2=500kPa, h_f=640.09 kJ\/kg, h_g=2748.01 kJ\/Kg"

The enthalpy of saturated steam "H_2 =m*h_2=mh_g=8*2748.01=21984.08 kJ"

If 70% of steam is condensed at state-3, the dryness fraction at state-3 is X=0.3.

Therefore enthalpy of steam at state -3 is given by

"H_3=m(hf+X(hg-hf))=10179.4 kJ"